A triangle has the following sides: 12 feet, 20 feet, and 16 feet. Is this a right triangle? Explain your answer.
Answer by edjones(8007)

If it is a right triangle then the hypotenuse=20 and:
a^2+b^2=c^2
12^2+16^2=400
20^2=200
It is a right triangle.
Ed